# What is the Plausible Memory Representation?

### v1

@@ Appendix F Comparision between TG and TG with a weak baseline regression model
| TG | 97.3 (4.0) | tg_nae | 2.93 ( 0.1) | 0.001 | 3/3 |
| tg_nae_so | 0.657 ( 0.1) | 0.01 | 3/3 | ||
| TG | 92.2 (10) | tg_mcomp_nae | 2.88 ( 0.2) | 0.001 | 3/3 |
| tg_mcomp_nae_so | 0.416 ( 0.1) | 0.01–0.17 | 1/3 |
In Section 5.1, we explored the advantage of treating closed phrases as single representations beyond explicit syntactic structure consideration. Our analysis incorporated Transformer’s NAE in the baseline regression model to distinguish between two effects in TG: syntactic structure consideration and direct terminal token access.
To evaluate which model—TG or TG—better captures more dominant factors in human sentence processing as a single model, we assessed their predictive power without Transformer’s NAE in the baseline regression model (Table 8). The analysis revealed TG’s superior predictive power (LogLik=97.3) compared to TG (LogLik=92.2). Additionally, TG demonstrated both immediate and spillover effects, while TG primarily showed an immediate effect. These results highlight that TG, which explicitly treats closed phrases as single representations, outperforms TG, even when considering TG’s advantage in direct terminal token access. Consistent with findings in Section 5.1, likelihood tests confirmed TG’s independent predictive power from TG (‘TG & TG’‘TG’, ); TG, despite its lower overall predictive power, accounted for unique variance (‘TG & TG’‘TG’, ).

### v2

@@ Appendix F Direct comparison of TG and TG
TG | 78.2 (6.9) | tg_nae | 1.43 ( 0.2) | 0.001 | 3/3 |
tg_nae_so | 2.28 ( 0.3) | 0.001 | 3/3 | ||
TG | 59.4 (17) | tg_mcomp_nae | 1.36 ( 0.2) | 0.001 | 3/3 |
tg_mcomp_nae_so | 1.90 ( 0.2) | 0.001 | 3/3 |
In Section 5.1, we explored the advantage of treating closed phrases as single representations beyond syntactic structure consideration. Our analysis incorporated Transformer’s NAE in the baseline regression model to distinguish between two effects in TG: syntactic structure consideration and direct terminal token access.
To evaluate which model—TG or TG—better captures more dominant factors in human sentence processing as a single model, we assessed their predictive power without Transformer’s NAE in the baseline regression model (Table 8). The analysis revealed TG’s superior predictive power (LogLik=78.2) compared to TG (LogLik=59.4). These results highlight that TG, which explicitly treats closed phrases as single representations, outperforms TG, even when considering TG’s advantage in direct terminal token access. The likelihood ratio tests confirmed TG’s independent predictive power from TG (‘TG & TG’‘TG’, ), while in contrast to Section 5.1, TG also accounted for unique variance (‘TG & TG’‘TG’, ). This bidirectional relationship likely emerges because TG’s direct terminal token access explains unique variance in the absence of Transformer’s NAE.